OpenStudy (anonymous):

please tell to how to solve question related to calclus and where &how to use integration &differentiation?

OpenStudy (anonymous):

rather a broad question, isn't it?

hartnn (hartnn):

Calculus Questions are easy to solve once you know the formulas :) integration is used in finding are of solids, figures, etc. differentiation can be used to find rate of change of a quantity. and both of these have many other applications.. if you have any specific question, you please ask :)

hartnn (hartnn):

finding area*

OpenStudy (anonymous):

\[gm/l \int\limits_{-l/2}^{+l/2}(r-x)^2dx\]

hartnn (hartnn):

you need to first expand (r-x)^2 =.... ?

OpenStudy (anonymous):

don't know

hartnn (hartnn):

\[(a-b)^2=a^2-2ab+b^2 \\ (r-x)^2 =... ?\]

OpenStudy (anonymous):

\[r^2-2rx+x^2\]

hartnn (hartnn):

thats correct, now \[\int x^n dx=x^{n+1}/(n+1) \\ \int x^2dx=... ?\]

OpenStudy (anonymous):

x^3/4

hartnn (hartnn):

it will be x^3/3 , as n+1 =3 \[ \int\limits_{-l/2}^{+l/2}(x^2-2rx+r^2)dx\\= \int\limits_{-l/2}^{+l/2}x^2dx-2r\int\limits_{-l/2}^{+l/2}x dx+r^2\int\limits_{-l/2}^{+l/2}1dx = \\ =[x^3/3]_{-l/2}^{+l/2} -2r [x^2/2]_{-l/2}^{+l/2}+r^2[x]_{-l/2}^{+l/2}\] got this ?

OpenStudy (anonymous):

why u hav taken 2r &r out?

hartnn (hartnn):

because when we are integrating with respect to x, 'r' will be constant and can be taken out of integration and differentiation..

OpenStudy (anonymous):

how u will know that x is a constant? or any digit or letter is a constant?

hartnn (hartnn):

do you see 'dx' in integration, ? means only 'x' is variable, all other letters are treated to be constant.

OpenStudy (anonymous):

the letter attach to d like dx always only that letter r variable?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

sorry, yes. only that letter is variable. so here r is constant, x is variable.

OpenStudy (anonymous):

okay after that

hartnn (hartnn):

now l/2 is upper limit and -l/2 is lower limit. \[[x^3/3]^{l/2}_{-l/2}\] is solved by first putting x= upper limit - x= lower limit like this : (l/2)^3/3- (-l/2)^3/3 got this first term ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

can you try for other 2 terms ?

hartnn (hartnn):

\[\int\limits_{-l/2}^{+l/2}(x^2-2rx+r^2)dx\\= \int\limits_{-l/2}^{+l/2}x^2dx-2r\int\limits_{-l/2}^{+l/2}x dx+r^2\int\limits_{-l/2}^{+l/2}1dx = \\ =[x^3/3]_{-l/2}^{+l/2} -2r [x^2/2]_{-l/2}^{+l/2}+r^2[x]_{-l/2}^{+l/2}= \\ = [(l/2)^3/(3)-(-l/2)^3/(3)]-2r[(l/2)^2/(2)-(-l/2)^2/(2)]\\ +r^2[l/2-(-l/2)] \\ = 2*l^3/24 -2r(2*l^2/8)+r^2(2l/2) \\ =l^3/12 - rl^2/2+r^2l\]

OpenStudy (anonymous):

hey this one is nt d answer.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

plz can u provide me conversion chats

OpenStudy (anonymous):

example : how to convert kg into mass. like this can u provide me the chat in which about all unit conversion given.